4. Find Φ12(x) over Q.



1
Expert's answer
2021-07-13T11:26:11-0400

Solution

We know that there are "\phi" (12) = 4 primitive 12th roots of unity. If

ω is one such a root, the others are "\omega^5" , "\omega^7" and "\omega^{11}". We know that

(x12− 1) = (x6 − 1)(x6+ 1) .

"\phi" (12) is a factor of (x6 + 1).

Furthermore, we know that

i and -i are non-primitive roots of unity, so we can remove a factor of x2 + 1, to get x4-x2+1.

Since it is a polynomial of 4 degree

It's the required one.

Answer:

"\phi_{12}(x)=x^4-x^2+1"